The perimeter of a rectangular Playing field is 504. It's length is 6 yards shorter than twice it's width. What is the area
Answer by ewatrrr(24785)


Hi,
length is 6 yards shorter than twice it's width
Let x and (2x-6) represent the width and length respectively
Question states*** P = 504
504 = x + x + (2x-6) +(2x-6)
solving for x
504 = 6x - 12
516 = 6x
86 = x, the width, Length is 166 (2*86 - 6)
CHECKING our Answer
2*86 + 2*166 = 172 + 332 = 504